# Prime Geodesic Theorem for Arithmetic Compact Surfaces

### v1

@@ 3. Stable Orbital Integrals
Let $\mathcal{B}$ be a quaternion algebra over $\mathbb{Q}$ with reduce trace ${\rm Tr}={\rm Tr}_{\mathcal{B}}$ and norm $\nu=\nu_{\mathcal{B}}$. Fix a maximal $\mathbb{Z}$-order $\mathcal{O}_{\mathcal{B}}$ of $\mathcal{B}$. Suppose we have two (congruence) subgroups $\Gamma\triangleleft\widetilde{\Gamma}<\mathcal{O}_{\mathcal{B}}^{\times}$, together with subgroups in finite adeles $\mathbf{K}_{\Gamma}\triangleleft\mathbf{K}_{\widetilde{\Gamma}}<\widehat{\mathcal{O}}_{\mathcal{B}}^{\times}\subset\mathcal{B}^{\times}(\mathbb{A}_{{\rm fin}})$, where $\mathbb{A}_{{\rm fin}}$ is the ring of finite adeles of $\mathbb{Q}$, such that
@@
Suppose $\mathcal{B}$ is unramified at $\infty$ from now on, namely $\mathcal{B}_{\infty}\simeq{\rm M}_{2}(\mathbb{R})$. Then $\Gamma^{1}$, resp. $\pm\Gamma^{1}$ can be regarded as a lattice in ${\rm SL}_{2}(\mathbb{R})$, resp. ${\rm PSL}_{2}(\mathbb{R})$.
@@
where $x_{0}=x_{0}([\gamma])=N(\gamma_{0})$ for a prmitive conjugacy class associated with $\gamma$. Note that the formula (3.2) is a group-theoretic reformulation of [6, Proposition I.6.3]. In particular, $\widetilde{g}_{\infty}$ is the Harish-Chandra transform of $\widetilde{f}_{\infty}$ and can be any even function in ${\rm C}_{c}^{\infty}(\mathbb{R})$ by [8, Proposition 3.4]. For any $t\in\mathbb{R}$ we define
(3.3) $\mathrm{d}\Psi_{\Gamma^{1}}(t):=\frac{1}{\sqrt{x}-\sqrt{x}^{-1}}\sum_{[\gamma]:{\rm Tr}(\gamma)=t}\log x_{0}([\gamma]).$
@@
(3.4) $\Psi_{\Gamma^{1}}(x)=c_{\Gamma^{1}}\cdot\sum_{2<\lvert t\rvert\leq x^{1/2}+x^{-1/2}}\mathrm{d}\Psi_{\Gamma^{1}}(t),\quad c_{\Gamma^{1}}:=\frac{[\pm\Gamma^{1}:\Gamma^{1}]}{2}=\begin{cases}1&\text{if }-\mathbbm{1}_{\mathcal{B}}\notin\Gamma^{1}\\ 1/2&\text{if }-\mathbbm{1}_{\mathcal{B}}\in\Gamma^{1}\end{cases}.$
@@
Suppose the group $\mathbf{K}_{\widetilde{\Gamma}}$ satisfies $\mathcal{B}^{\times}(\mathbb{A}_{{\rm fin}})=\mathcal{B}^{\times}(\mathbb{Q})\cdot\mathbf{K}_{\widetilde{\Gamma}}$. Let $\gamma\in\Gamma^{1}$ be hyperbolic with ${\rm Tr}(\gamma)=t$ (hence $\lvert t\rvert>2$).
(1) Such $\gamma$ is $\mathbb{Q}$-elliptic, i.e., the $\mathbb{Q}$-algebra $\mathbb{Q}[\gamma]$ is a (real) quadratic field.
@@ 4. Proof of Main Theorem
(4.1) $c_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\cdot\widetilde{g}_{\infty}(\log x)\cdot\mathrm{d}\Psi_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\left(t\right)=\\ {\rm Vol}(\widehat{\mathbb{Z}}^{\times}){\rm Vol}\left(\mathbf{Z}_{\gamma}(\mathbb{Q})\mathbb{A}^{\times}\backslash\mathbf{Z}_{\gamma}(\mathbb{A})\right)\int_{\mathbf{Z}_{\gamma}(\mathbb{A})\backslash\mathcal{D}^{\times}(\mathbb{A})}\varphi_{\mathcal{D}}(g^{-1}\gamma g)\mathrm{d}g.$
@@
For each $I\subset{\rm Ram}(\mathcal{D})$, let $\mathcal{O}_{I}$ be the Eichler order, whose completion $\mathcal{O}_{I,p}$ at $p\in I$ and $p\notin{\rm Ram}(\mathcal{D})$ is $\mathfrak{M}_{p}:={\rm M}_{2}(\mathbb{Z}_{p})$ and at $p\in{\rm Ram}(\mathcal{D})-I$ is $\mathfrak{J}_{p}:=\begin{pmatrix}\mathbb{Z}_{p}&\mathbb{Z}_{p}\\ p\mathbb{Z}_{p}&\mathbb{Z}_{p}\end{pmatrix}$. Let $\mathfrak{N}_{I}$ be the two sided ideal of $\mathcal{O}_{I}$, whose completion at $p\in I$, resp. $p\notin{\rm Ram}(\mathcal{D})$, resp. $p\in{\rm Ram}(\mathcal{D})-I$ is the $n_{p}^{1}$-th, resp. $n_{p}$-th, resp. $n_{p}^{2}$-th power of the Jacobson radical of $\mathcal{O}_{I,p}$. Define
@@
It follows from Proposition 3.1 again that
(4.2) $c_{\Gamma^{1}(\mathfrak{N}_{I})}\cdot\widetilde{g}_{\infty}(\log x)\cdot\mathrm{d}\Psi_{\Gamma^{1}(\mathfrak{N}_{I})}\left(t\right)={\rm Vol}(\widehat{\mathbb{Z}}^{\times}){\rm Vol}\left(\mathbf{Z}_{\gamma}(\mathbb{Q})\mathbb{A}^{\times}\backslash\mathbf{Z}_{\gamma}(\mathbb{A})\right)\int_{\mathbf{Z}_{\gamma}(\mathbb{A})\backslash{\rm GL}_{2}(\mathbb{A})}f_{I}(g^{-1}\gamma g)\mathrm{d}g.$
@@
$c_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\cdot\widetilde{g}_{\infty}(\log x)\cdot\mathrm{d}\Psi_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\left(t\right)=\sum_{I\subset{\rm Ram}(\mathcal{D})}\left(\prod_{p\in I}a_{p}\prod_{p\in{\rm Ram}(\mathcal{D})-I}b_{p}\right)c_{\Gamma^{1}(\mathfrak{N}_{I})}\cdot\widetilde{g}_{\infty}(\log x)\cdot\mathrm{d}\Psi_{\Gamma^{1}(\mathfrak{N}_{I})}\left(t\right),$
@@
(4.4) $c_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\cdot\mathrm{d}\Psi_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\left(t\right)=\sum_{I\subset{\rm Ram}(\mathcal{D})}\left(\prod_{p\in I}a_{p}\prod_{p\in{\rm Ram}(\mathcal{D})-I}b_{p}\right)c_{\Gamma^{1}(\mathfrak{N}_{I})}\cdot\mathrm{d}\Psi_{\Gamma^{1}(\mathfrak{N}_{I})}\left(t\right).$
@@
Now that the prime geodesic theorem hold for $\Gamma(\mathfrak{N}_{I})$ by [10, Corollary 1.2] in the form
@@
we conclude the proof of Theorem 1.1 by the crucial identity
@@
since we have $a_{p}+b_{p}=1$ by observing Proposition 2.4.

### v2

@@ 3. Stable Orbital Integrals
Let $\mathcal{B}$ be a quaternion algebra over $\mathbb{Q}$ with reduced trace ${\rm Tr}={\rm Tr}_{\mathcal{B}}$ and norm $\nu=\nu_{\mathcal{B}}$. Suppose $\mathcal{B}$ is unramified at $\infty$. Fix a maximal $\mathbb{Z}$-order $\mathcal{O}_{\mathcal{B}}$ of $\mathcal{B}$. Suppose we have two (congruence) subgroups $\Gamma\triangleleft\widetilde{\Gamma}<\mathcal{O}_{\mathcal{B}}^{\times}$, together with subgroups in finite adeles $\mathbf{K}_{\Gamma}\triangleleft\mathbf{K}_{\widetilde{\Gamma}}<\widehat{\mathcal{O}}_{\mathcal{B}}^{\times}\subset\mathcal{B}^{\times}(\mathbb{A}_{{\rm fin}})$, where $\mathbb{A}_{{\rm fin}}$ is the ring of finite adeles of $\mathbb{Q}$, so that we have
$\Gamma=\mathcal{B}^{\times}(\mathbb{Q})\cap\mathbf{K}_{\Gamma},\quad\widetilde{\Gamma}=\mathcal{B}^{\times}(\mathbb{Q})\cap\mathbf{K}_{\widetilde{\Gamma}},\quad\mathbf{K}_{\widetilde{\Gamma}}\cap\mathbb{A}_{{\rm fin}}^{\times}=\widehat{\mathbb{Z}}^{\times}$
and the following crucial assumption as well as a consequence of it via [13, Theorem 28.5.3] (the strong approximation theorem)
(3.1) $\nu(\mathbf{K}_{\widetilde{\Gamma}})=\widehat{\mathbb{Z}}^{\times}\quad\Rightarrow\quad\mathcal{B}^{\times}(\mathbb{A}_{{\rm fin}})=\mathcal{B}^{\times}(\mathbb{Q})\cdot\mathbf{K}_{\widetilde{\Gamma}}$.
Remark 3.1.
For example for $\mathcal{B}={\rm M}_{2}$, the following groups satisfy the above assumptions ($N\in\mathbb{Z}_{\geq 1}$):
  • (1)
@@
Since $\mathcal{B}_{\infty}\simeq{\rm M}_{2}(\mathbb{R})$, the discrete group $\Gamma^{1}$, resp. $\pm\Gamma^{1}/\{\pm\mathbbm{1}\}$ can be regarded as a lattice in ${\rm SL}_{2}(\mathbb{R})$, resp. ${\rm PSL}_{2}(\mathbb{R})={\rm SL}_{2}(\mathbb{R})/\{\pm\mathbbm{1}\}$.
@@
where $x_{0}=x_{0}([\gamma])=N(\gamma_{0})$ for a primitive conjugacy class associated with $\gamma$. Note that the formula (3.2) is a group-theoretic reformulation of [6, Proposition I.6.3]. In particular, $\widetilde{g}_{\infty}$ is the Harish-Chandra transform of $\widetilde{f}_{\infty}$ and can be any even function in ${\rm C}_{c}^{\infty}(\mathbb{R})$ by [8, Proposition 3.4]. For any $t\in\mathbb{R}$ we define
(3.3) $\psi_{\Gamma^{1}}(t):=\frac{1}{1-x^{-1}}\sum_{[\gamma]:{\rm Tr}(\gamma)=t}\log x_{0}([\gamma]).$
@@
(3.4) $\Psi_{\Gamma^{1}}(x)=c_{\Gamma^{1}}\cdot\sum_{2<\lvert t\rvert\leq x^{1/2}+x^{-1/2}}\psi_{\Gamma^{1}}(t),\quad c_{\Gamma^{1}}:=\frac{[\pm\Gamma^{1}:\Gamma^{1}]}{2}=\begin{cases}1&\text{if }-\mathbbm{1}_{\mathcal{B}}\notin\Gamma^{1}\\ 1/2&\text{if }-\mathbbm{1}_{\mathcal{B}}\in\Gamma^{1}\end{cases}.$
@@
(0) The crucial assumption (3.1) implies $\mathcal{B}^{\times}(\mathbb{A}_{{\rm fin}})=\mathcal{B}^{\times}(\mathbb{Q})\cdot\mathbf{K}_{\widetilde{\Gamma}}$.
(1) Let $\gamma\in\Gamma^{1}$ be hyperbolic with ${\rm Tr}(\gamma)=t$ (hence $\lvert t\rvert>2$). Such $\gamma$ is $\mathbb{Q}$-elliptic, i.e., the $\mathbb{Q}$-algebra $\mathbb{Q}[\gamma]$ is a (real) quadratic field.
@@
(0) This follows directly from the strong approximation theorem [13, Theorem 28.5.3].
@@ 4. Proof of Main Theorem
(4.1) $c_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\cdot\frac{\widetilde{g}_{\infty}(\log x)}{\sqrt{x}}\cdot\psi_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\left(t\right)=\\ {\rm Vol}(\widehat{\mathbb{Z}}^{\times}){\rm Vol}\left(\mathbf{Z}_{\gamma}(\mathbb{Q})\mathbb{A}^{\times}\backslash\mathbf{Z}_{\gamma}(\mathbb{A})\right)\int_{\mathbf{Z}_{\gamma}(\mathbb{A})\backslash\mathcal{D}^{\times}(\mathbb{A})}\varphi_{\mathcal{D}}(g^{-1}\gamma g)\mathrm{d}g.$
@@
For each $I\subset{\rm Ram}(\mathcal{D})$, let $\mathcal{O}_{I}$ be the Eichler order, whose completion $\mathcal{O}_{I,p}$ at $p\in I$ is $\mathfrak{M}_{p}:={\rm M}_{2}(\mathbb{Z}_{p})$ and at $p\in{\rm Ram}(\mathcal{D})-I$ is $\mathfrak{J}_{p}:=\begin{pmatrix}\mathbb{Z}_{p}&\mathbb{Z}_{p}\\ p\mathbb{Z}_{p}&\mathbb{Z}_{p}\end{pmatrix}$. Let $\mathfrak{N}_{I}$ be the two sided ideal of $\mathcal{O}_{I}$, whose completion at $p\in I$, resp. $p\notin{\rm Ram}(\mathcal{D})$, resp. $p\in{\rm Ram}(\mathcal{D})-I$ is the $n_{p}^{1}$-th, resp. $n_{p}$-th, resp. $n_{p}^{2}$-th power of the Jacobson radical of $\mathcal{O}_{I,p}$. Define
@@
It follows from Proposition 3.2 (taking into account Remark 3.1 (3)) again that
(4.2) $c_{\Gamma^{1}(\mathfrak{N}_{I})}\cdot\frac{\widetilde{g}_{\infty}(\log x)}{\sqrt{x}}\cdot\psi_{\Gamma^{1}(\mathfrak{N}_{I})}\left(t\right)={\rm Vol}(\widehat{\mathbb{Z}}^{\times}){\rm Vol}\left(\mathbf{Z}_{\gamma}(\mathbb{Q})\mathbb{A}^{\times}\backslash\mathbf{Z}_{\gamma}(\mathbb{A})\right)\int_{\mathbf{Z}_{\gamma}(\mathbb{A})\backslash{\rm GL}_{2}(\mathbb{A})}f_{I}(g^{-1}\gamma g)\mathrm{d}g.$
@@
$c_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\cdot\frac{\widetilde{g}_{\infty}(\log x)}{\sqrt{x}}\cdot\psi_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\left(t\right)=\sum_{I\subset{\rm Ram}(\mathcal{D})}\left(\prod_{p\in I}a_{p}\prod_{p\in{\rm Ram}(\mathcal{D})-I}b_{p}\right)c_{\Gamma^{1}(\mathfrak{N}_{I})}\cdot\frac{\widetilde{g}_{\infty}(\log x)}{\sqrt{x}}\cdot\psi_{\Gamma^{1}(\mathfrak{N}_{I})}\left(t\right),$
@@
(4.4) $c_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\cdot\psi_{\Gamma_{\mathcal{D}}^{1}(\mathfrak{N})}\left(t\right)=\sum_{I\subset{\rm Ram}(\mathcal{D})}\left(\prod_{p\in I}a_{p}\prod_{p\in{\rm Ram}(\mathcal{D})-I}b_{p}\right)c_{\Gamma^{1}(\mathfrak{N}_{I})}\cdot\psi_{\Gamma^{1}(\mathfrak{N}_{I})}\left(t\right).$
@@
Now that the prime geodesic theorem holds for $\Gamma(\mathfrak{N}_{I})$ by [10, Corollary 1.2] in the form
@@
we conclude the proof of Theorem 1.1 for $\Psi_{\Gamma_{\mathcal{D}}(\mathfrak{N})}$ by the crucial identity
@@
since we have $a_{p}+b_{p}=1$ by observing Proposition 2.5. Finally to derive the statement for $\pi_{\Gamma_{\mathcal{D}}(\mathfrak{N})}$ we simply do an integration by parts (or abel summation).